# Track4Gen: Teaching Video Diffusion Models to Track Points Improves Video Generation

### v2

@@ Abstract
While recent foundational video generators produce visually rich output, they still struggle with appearance drift, where objects gradually degrade or change inconsistently across frames, breaking visual coherence. We hypothesize that this is because there is no explicit supervision in terms of spatial tracking at the feature level. We propose Track4Gen, a spatially aware video generator that combines video diffusion loss with point tracking across frames, providing enhanced spatial supervision on the diffusion features. Track4Gen merges the video generation and point tracking tasks into a single network by making minimal changes to existing video generation architectures. Using Stable Video Diffusion [5] as a backbone, Track4Gen demonstrates that it is possible to unify video generation and point tracking, which are typically handled as separate tasks. Our extensive evaluations show that Track4Gen effectively reduces appearance drift, resulting in temporally stable and visually coherent video generation. Project page: hyeonho99.github.io/track4gen
footnotetext: Work mostly done while Hyeonho Jeong was an intern at Adobe.
@@ 1 Introduction
We train Track4Gen using the latest Stable Video Diffusion [5] as the backbone and evaluate on the publicly available VBench dataset [32, 33]. We report significant improvement in terms of appearance constancy of subjects, both in quantitative and qualitative (i.e., via user studies) evaluations. In summary, we demonstrate that it is possible to upgrade existing video generators, by supervising them with additional correspondence tracking loss, to produce videos without significant appearance drifts, a problem commonly encountered in diffusion-based video generators.
@@ 2 Related Work
With advances in generation, systematic evaluation of generation quality has become crucial. Traditionally, metrics such as Fréchet Inception Distance (FID, [28]), Fréchet Video Distance (FVD, [68]), and CLIPSIM [54] are used. Additionally, comprehensive benchmark suites [32, 33, 73] have been introduced to provide a more robust evaluation aligned with human perception. Inspired by such work, we thoroughly evaluate our approach and demonstrate improved video generation quality with respect to both conventional metrics and the recent VBench metrics [32].
@@
Various foundational models such as vision transformers [17] or diffusion-based generators [55] have been utilized as feature extractors for various tasks including semantic matching [27, 46, 18], classification [43], segmentation [76, 71], and editing [66, 21, 23, 36]. There have been efforts to boost their performance by post-processing the feature maps obtained from the pre-trained models, e.g., by upsampling [20, 63]. In a recent effort, Yue et al. [80] lift semantic per-frame features from a foundational model into a 3D Gaussian representation. They fine-tune the foundational model with such 3D-aware features resulting in improved performance in downstream tasks. Similarly, Sundaram et al. [62] fine-tune state-of-the-art foundational models on human similarity judgments yielding improved representations across downstream tasks. In a concurrent effort, Yu et al. [79] propose to align the internal features of an image generation model with external discriminative features [50], which results in more effective training of the generator.
@@ 3.1 Background: Stable Video Diffusion
Figure 3: Real-world video tracking using different video diffusion features. Given color-coded query points on the first frame (Leftmost column), we display tracked points on target frames using features from different blocks (right columns). The 13th frame (first row) and 8th frame (second row) are shown as target frames. Full results are available on our page.
@@ 3.2 Video Diffusion Features
Figure 5: Image-to-video generation results of the original SVD and Track4Gen. Please visit our project for full video view.
@@ 4.1 Implementation Details
To train Track4Gen, we construct a training dataset consisting of $567$ video-trajectory pairs, with each video having a resolution of $320\times 576$ and a duration of $24$ frames. Since no real-world video with (dense) ground-truth trajectory annotations exist at the time of this work, we utilize optical flow to generate trajectory annotations. For details, see Sec. A.1. A key challenge is the need for accurate video segmentation maps to ensure a balanced distribution of trajectory points between foreground objects and the background [14]. To address this, we utilize public video datasets paired with ground-truth segmentation maps [44, 51, 53, 8, 19], where we split longer videos into 24-frame segments.
@@ 4.2 Track4Gen for Video Generation
Automatic metrics. We first report five key metrics from VBench [32]: (1) Subject Consistency—assesses subject appearance consistency of the video by computing the similarity of DINO [50] features. (2) Temporal Flickering—detects temporal consistency by taking static frames and calculating the mean absolute difference across frames. (3) Motion Smoothness—measures smoothness of motion, and how well it adheres to real-world physics, using video frame interpolation priors [47]. (4) Image Quality—evaluates distortions (e.g., noise, blur) using a pretrained, multi-scale image quality predictor [40]. (5) Video-Image Alignment—measures alignment between the subject in the input image and in the generated video using DINO features. We additionally report FID [28] and FVD [68]. To further assess temporal consistency, CLIPSIM [54] and LPIPS [84] are reported in the Appendix (see Sec. B).
@@
Human evaluation. We further evaluate Track4Gen against baselines through a user study. We ask $64$ participants to compare our results with a randomly selected baseline. We ask the users to evaluate how consistent main objects appear across the frames in a generated video as well as how natural the depicted motion is. We provide further details of the user study in Sec. A.3 of the Appendix.
@@ 4.3.1 Zero-shot Feature Comparison
Datasets. We use TAP-Vid DAVIS [14] and BADJA [4] as benchmark datasets. Additionally, we include two shorter benchmarks, DAVIS (24-frame) and BADJA (24-frame), which focus on the first 24 frames with query and target points within this range. Details on encoding long videos using pretrained video diffusion models are in Sec. A.4 of the Appendix.
@@ 4.3.2 Extending Track4Gen with Test-time Adaptation
Tab. 3 shows that Track4Gen features optimized with [67] achieve performance comparable to dedicated trackers. Qualitative results are in Fig. 9 and in Fig. 15, 17. In addition, the adaptation progress is visualized in Fig. 17.
@@ 4.4 Ablation Studies
We present an ablation study in Tab. 4 where we train different set of modules. Each spatio-temporal block of SVD includes both spatial and temporal transformers. We compare training only spatial transformers, only temporal transformers, or both. We also ablate the architecture of the refiner module using either 2D or 3D convolution layers. Our analysis shows that while results are similar across settings, training only the temporal transformers in SVD with 2D convolutions as the refiner module yields optimal video generation quality. We further analyze our training dataset by additionally incorporating Kubric [22] simulated videos (1K video-tracks pairs from the Panning MOVi-E data [15, 12]) with automatically annotated trajectories into training. As shown in Tab. 5, optical flow-chained tracklets from real videos provide as effective correspondence guidance as tracklets from synthetic data, while synthetic videos negatively impact the video generation quality.
@@ 5 Conclusion and Future Work
We have presented the first unified framework that bridges two distinct tasks: video generation and dense point tracking. We demonstrated that this produces temporally consistent feature representations and appearance-consistent videos. As for limitations, videos generated by Track4Gen tend to exhibit less dynamic motion compared to those from other video generators. Additionally, failure cases are included in the Appendix.
Future work. Recently, cutting-edge video trackers [12, 16, 38] have emerged, enabling dense, accurate, and long-term tracking, especially with better handling of occlusions. This opens up promising future directions for extending our work to utilize real-world videos, automatically annotated by these advanced trackers. Additionally, we aim to explore conditional video generation, using point tracks to guide or blend motions as a coarse authoring tool.
Acknowledgments. We would like to thank Seokju Cho and Narek Tumanyan for their invaluable feedback on video point tracking. We also extend our gratitude to Mingi Kwon, Joon-Young Lee, and Gabriel Huang for their insightful remarks and discussions.
Appendix
This document is structured as follows: Sec. A provides additional implementation details for the experiments. In Sec. B, we report supplementary quantitative metrics for video generation assessment. Sec. C presents additional qualitative results for image-to-video generation, while Sec. D focuses on qualitative video tracking results. Following this, we discuss the potential limitations and failure cases of Track4Gen in Sec. E.
@@ A.1 Preprocessing Video Correspondence
We utilize RAFT optical flow [65] to compute dense point trajectories across video frames. RAFT has demonstrated robust point tracking performance across various input types [70], even compared to supervised trackers like TAP-Net [14]. Following previous tracking literature [70, 67], we first compute pairwise correspondences between all consecutive frames. Tracks are then formed by chaining the estimated flow fields and filtered using a cycle consistency constraint. Specifically, given a point $\boldsymbol{\text{x}}^{i}$ in frame $i$ and optical flow between frames $i$ and $i+1$ denoted as $\boldsymbol{f}_{i\rightarrow i+1}$, the corresponding point in frame $i+1$ is estimated as $\boldsymbol{\text{x}}^{i+1}=\boldsymbol{\text{x}}^{i}+\boldsymbol{f}_{i\rightarrow i+1}(\boldsymbol{\text{x}}^{i})$. We retain the pair $(\boldsymbol{\text{x}}^{i},\boldsymbol{\text{x}}^{i+1})$ only if it satisfies $\bigl{|}\bigl{|}\boldsymbol{\text{x}}^{i}-(\boldsymbol{\text{x}}^{i+1}+\boldsymbol{f}_{i+1\rightarrow i}(\boldsymbol{\text{x}}^{i+1}))\bigl{|}\bigl{|}\leq 1.5$.
@@ C.1 Comparisons
In Fig. 13 and 14, we present a comparison of Track4Gen against all three baselines: (1) the pretrained Stable Video Diffusion, (2) Stable Video Diffusion finetuned without the tracking loss, and (3) Track4Gen trained without the Refiner module. For a better view, please visit this page.
@@ D.2 Track4Gen with DINO-Tracker
Appendix E Discussion on Limitation and Failure
For video results related to this section, please visit this page and the bottom of this page. While Track4Gen significantly enhances appearance constancy in generated videos, it tends to result in reduced camera motion compared to the original Stable Video Diffusion prior, a behavior also observed in the finetuned Stable Video Diffusion baseline. (see Fig. 20). We attribute this to the training dataset used for finetuning. In addition, in some cases Track4Gen produces unrealistic motion and exhibit artifacts on human faces and hands, particularly when the resolution or size of the human subject in the video is small — a common limitation shared by video diffusion models [42], including the baselines. Typical failure cases of video generation are illustrated in Fig. 21.

### v3

@@ Abstract
While recent foundational video generators produce visually rich output, they still struggle with appearance drift, where objects gradually degrade or change inconsistently across frames, breaking visual coherence. We hypothesize that this is because there is no explicit supervision in terms of spatial tracking at the feature level. We propose Track4Gen, a spatially aware video generator that combines video diffusion loss with point tracking across frames, providing enhanced spatial supervision on the diffusion features. Track4Gen merges the video generation and point tracking tasks into a single network by making minimal changes to existing video generation architectures. Using Stable Video Diffusion [5] as a backbone, Track4Gen demonstrates that it is possible to unify video generation and point tracking, which are typically handled as separate tasks. Our extensive evaluations show that Track4Gen effectively reduces appearance drift, resulting in temporally stable and visually coherent video generation. Project page: hyeonho99.github.io/track4gen.
footnotetext: Work done during internship at Adobe.
@@ 1 Introduction
We train Track4Gen using the latest Stable Video Diffusion [5] as the backbone and evaluate on the publicly available VBench dataset [33]. We report significant improvement in terms of appearance constancy of subjects, both in quantitative and qualitative (i.e., via user studies) evaluations. In summary, we demonstrate that it is possible to upgrade existing video generators, by supervising them with additional correspondence tracking loss, to produce videos without significant appearance drifts, a problem commonly encountered in diffusion-based video generators.
@@ 2 Related Work
With advances in generation, systematic evaluation of generation quality has become crucial. Traditionally, metrics such as Fréchet Inception Distance (FID, [28]), Fréchet Video Distance (FVD, [67]), and CLIPSIM [53] are used. Additionally, comprehensive benchmark suites [33, 72] have been introduced to provide a more robust evaluation aligned with human perception. Inspired by such work, we thoroughly evaluate our approach and demonstrate improved video generation quality with respect to both conventional metrics and the recent VBench metrics [33].
@@
Various foundational models such as vision transformers [17] or diffusion-based generators [54] have been utilized as feature extractors for various tasks including semantic matching [27, 45, 18], classification [42], segmentation [75, 70], and editing [65, 21, 23]. There have been efforts to boost their performance by post-processing the feature maps obtained from the pre-trained models, e.g., by upsampling [20, 62]. In a recent effort, Yue et al. [79] lift semantic per-frame features from a foundational model into a 3D Gaussian representation. They fine-tune the foundational model with such 3D-aware features resulting in improved performance in downstream tasks. Similarly, Sundaram et al. [61] fine-tune state-of-the-art foundational models on human similarity judgments yielding improved representations across downstream tasks. In a concurrent effort, Yu et al. [78] propose to align the internal features of an image generation model with external discriminative features [49], which results in more effective training of the generator.
@@ 3.1 Background: Stable Video Diffusion
Figure 3: Real-world video tracking using different video diffusion features. Given color-coded query points on the first frame (Leftmost column), we display tracked points on target frames using features from different blocks (right columns). The 13th frame (first row) and 8th frame (second row) are shown as target frames. Full results are available in the supplementary and on our page.
@@ 3.2 Video Diffusion Features
Figure 5: Image-to-video generation results of the original SVD and Track4Gen. Please visit our page for full video view.
@@ 4.1 Implementation Details
To train Track4Gen, we construct a training dataset consisting of $567$ video-trajectory pairs, with each video having a resolution of $320\times 576$ and a duration of $24$ frames. Since no real-world video with (dense) ground-truth trajectory annotations exist at the time of this work, we utilize optical flow to generate trajectory annotations. A key challenge is the need for accurate video segmentation maps to ensure a balanced distribution of trajectory points between foreground objects and the background [14]. To address this, we utilize public video datasets paired with ground-truth segmentation maps [43, 50, 52, 8, 19], where we split longer videos into 24-frame segments.
@@ 4.2 Track4Gen for Video Generation
Automatic metrics. We first report five key metrics from VBench [33]: (1) Subject Consistency—assesses subject appearance consistency of the video by computing the similarity of DINO [49] features. (2) Temporal Flickering—detects temporal consistency by taking static frames and calculating the mean absolute difference across frames. (3) Motion Smoothness—measures smoothness of motion, and how well it adheres to real-world physics, using video frame interpolation priors [46]. (4) Image Quality—evaluates distortions (e.g., noise, blur) using a pretrained, multi-scale image quality predictor [39]. (5) Video-Image Alignment—measures alignment between the subject in the input image and in the generated video using DINO features. We additionally report FID [28] and FVD [67].
@@
Human evaluation. We further evaluate Track4Gen against baselines through a user study. We ask $64$ participants to compare our results with a randomly selected baseline. We ask the users to evaluate how consistent main objects appear across the frames in a generated video as well as how natural the depicted motion is. We provide further details of the user study in the supplementary material.
@@ 4.3.1 Zero-shot Feature Comparison
Datasets. We use TAP-Vid DAVIS [14] and BADJA [4] as benchmark datasets. Additionally, we include two shorter benchmarks, DAVIS (24-frame) and BADJA (24-frame), which focus on the first 24 frames with query and target points within this range. Details on encoding long videos with the video models are in the supplemental.
@@ 4.3.2 Extending Track4Gen with Test-time Adaptation
Tab. 3 shows that Track4Gen features optimized with [66] achieve performance comparable to dedicated trackers. Qualitative results are in Fig. 9 and in the supplemental.
@@ 4.4 Ablation Studies
We present an ablation study in Tab. 4 where we train different set of modules. Each spatio-temporal block of SVD includes both spatial and temporal transformers. We compare training only spatial transformers, only temporal transformers, or both. We also ablate the architecture of the refiner module using either 2D or 3D convolution layers. Our analysis shows that while results are similar across settings, training only the temporal transformers in SVD with 2D convolutions as the refiner module yields optimal video generation quality. We further analyze our training dataset by additionally incorporating Kubric [22] simulated videos (1K video-track pairs from the Panning MOVi-E data [15, 12]) with automatically annotated trajectories into training. As shown in Tab. 5, optical flow-chained tracklets from real provides provide as effective correspondence guidance as tracklets from synthetic data, while synthetic videos negatively impact the video generation quality.
@@ 5 Conclusion and Future Work
We have presented the first unified framework that bridges two distinct tasks: video generation and dense point tracking. We demonstrated that this produces temporally consistent feature representations and appearance-consistent videos. As for limitations, videos generated by Track4Gen tend to exhibit less dynamic motion compared to those from other video generators. Additionally, failure cases are included in the supplementary material.
Future work. Recently, cutting-edge video trackers [12, 16, 37] have emerged, enabling dense, accurate, and long-term tracking, especially with better handling of occlusions. This opens up promising future directions for extending our work to utilize real-world videos, automatically annotated by these trackers to produce 3D-consistent videos [32].
Acknowledgments. We thank Seokju Cho and Narek Tumanyan for their invaluable feedback on video point tracking. We also extend our gratitude to Mingi Kwon, Joon-Young Lee, and Gabriel Huang for their insightful discussions. Hyeonho Jeong and Jong Chul Ye are supported by the National Research Foundation of Korea (NRF) under Grants RS-2024-00336454 and RS-2023-00262527, and by the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program, KAIST). Niloy J. Mitra is supported by UCL AI Centre.
Supplementary Material
This supplementary material is structured as follows: Sec. A provides additional implementation details for the experiments. In Sec. B, we report supplementary quantitative metrics for video generation assessment. Sec. C presents additional qualitative results for image-to-video generation, while Sec. D focuses on qualitative video tracking results. Following this, we discuss the potential limitations and failure cases of Track4Gen in Sec. E.
@@ A.1 Preprocessing Video Correspondence
We utilize RAFT optical flow [64] to compute dense point trajectories across video frames. RAFT has demonstrated robust point tracking performance across various input types [69], even compared to supervised trackers like TAP-Net [14]. Following previous tracking literature [69, 66], we first compute pairwise correspondences between all consecutive frames. Tracks are then formed by chaining the estimated flow fields and filtered using a cycle consistency constraint. Specifically, given a point $\boldsymbol{\text{x}}^{i}$ in frame $i$ and optical flow between frames $i$ and $i+1$ denoted as $\boldsymbol{f}_{i\rightarrow i+1}$, the corresponding point in frame $i+1$ is estimated as $\boldsymbol{\text{x}}^{i+1}=\boldsymbol{\text{x}}^{i}+\boldsymbol{f}_{i\rightarrow i+1}(\boldsymbol{\text{x}}^{i})$. We retain the pair $(\boldsymbol{\text{x}}^{i},\boldsymbol{\text{x}}^{i+1})$ only if it satisfies $\|\boldsymbol{\text{x}}^{i}-(\boldsymbol{\text{x}}^{i+1}+\boldsymbol{f}_{i+1\rightarrow i}(\boldsymbol{\text{x}}^{i+1}))\|_{2}\leq 1.5$, where $h{\times}w$ is set as $320{\times}576$. Also, a pair $(\boldsymbol{\text{x}}^{i},\boldsymbol{\text{x}}^{j})$ is filtered out if $\|\boldsymbol{\text{x}}^{j}{-}\boldsymbol{\text{x}}^{i\rightarrow j}\|_{2}{\geq 2}$ and $\|\boldsymbol{\text{x}}^{i}{-}(\boldsymbol{\text{x}}^{i\rightarrow j}{+}\boldsymbol{f}_{j\rightarrow i}(\boldsymbol{\text{x}}^{i\rightarrow j}))\|_{2}{\leq 1.5}$.
@@ C.1 Comparisons
In Fig. 13 and 14, we present a comparison of Track4Gen against all three baselines: (1) the pretrained Stable Video Diffusion, (2) Stable Video Diffusion finetuned without the tracking loss, and (3) Track4Gen trained without the Refiner module. For a better view, please visit page 2 (top) of our project page.
@@ D.2 Track4Gen with DINO-Tracker
Appendix E Discussion on Limitation and Failure
For video results related to this section, please refer to page 4 of our project page. While Track4Gen significantly enhances appearance constancy in generated videos, it tends to result in reduced camera motion compared to the original Stable Video Diffusion prior, a behavior also observed in the finetuned Stable Video Diffusion baseline. (see Fig. 20). We attribute this to the training dataset used for finetuning. In addition, in some cases Track4Gen produces unrealistic motion and exhibit artifacts on human faces and hands, particularly when the resolution or size of the human subject in the video is small — a common limitation shared by video diffusion models [41], including the baselines. Typical failure cases of video generation are illustrated in Fig. 21.